OpenStudy (rock_mit182):

Find an algorithm to determine if a number is a perfect number PLease help!

OpenStudy (rock_mit182):

\[F = Kx\]

OpenStudy (rock_mit182):

\[M *\frac{ L }{ T^2 } = K*L\]

OpenStudy (rock_mit182):

\[K = \frac{ M }{ L }* \frac{ L }{ T^2 }\]

OpenStudy (rock_mit182):

\[K = \frac{ M }{ T^2 }\]

OpenStudy (rock_mit182):

\[T = 2*\pi \sqrt{\frac{ m }{ k}}\]

OpenStudy (rock_mit182):

\[T = \sqrt{M} *\frac{ \sqrt{T^2} }{ \sqrt{M} }\]

OpenStudy (rock_mit182):

T = T

OpenStudy (rock_mit182):

\[T(s) \alpha \sqrt{m(g^\frac{ 1 }{ 2 }})\]